# Trustworthy AI on Safety, Bias, and Privacy: A Survey

### v1

@@ 2 Safety Alignment in LLMs
In Large Language Models, alignment aims to teach models human-desired behaviors and remove undesired behaviors. Safety alignment has often been treated as a subset of broader alignment challenges, with a primary focus on safety Li and Kim (2024). In this context, the goal of safety alignment is preventing LLMs from generating toxic or harmful content and simultaneously considers security problems in adversarial scenarios, such as jailbreak attempts Qi et al. (2024).
@@ 2.1 Why is Safety in LLMs Important?
With the release of AI services such as ChatGPT, Claude, and Gemini, AI-powered applications have become increasingly integrated into our daily lives, spanning fields like healthcare, finance, education, transportation, and even military applications OpenAI (2022); Gemini (2023). However, this rapid adoption also raises serious concerns regarding AI misuse. For instance, in 2023, the first suspected case of AI-assisted suicide was reported Brussels Times (2023), and AI-generated misinformation has been widely disseminated online, potentially manipulating public opinion Monteith et al. (2024). These incidents compel us to critically examine how to prevent AI from being misused in ways that harm society.
This challenge has become even more pressing with the rise of open-sourced large language models such as Llama and Deepseek families Touvron et al. (2023); Dubey et al. (2024); Guo et al. (2025), which allow individuals to control and finetune LLMs directly. As a result, the risks associated with AI misuse are expanding exponentially while government regulatory measures struggle to keep pace. Given these developments, AI safety has become a pressing need in the current research community.
@@ 2.2 How to Implement Safety in LLMs?
Upon the release of GPT-3 in 2020, we witnessed LLM’s remarkable language generation capabilities. However, concerns regarding bias, toxic content, and hallucinations also emerged, indicating that the model was still not ready for public release Brown et al. (2020). Later, in 2021, Anthropic introduced the HHH principle—Helpful, Honest, and Harmless—as a key principle of a truly beneficial AI assistant Askell et al. (2021). This concept set a foundational standard for AI assistants, clearly indicating that safety is an important objective of LLMs.
@@
The launch of ChatGPT in late 2022 marked a turning point, as it was the first large-scale exposure of LLMs to the public, allowing people to experience an AI assistant that felt genuinely helpful and capable of human-like reasoning. This breakthrough was largely built upon the techniques outlined in the InstructGPT, which introduced Supervised Fine-Tuning (SFT) and Reinforcement Learning with Human Feedback (RLHF) as key methods for aligning LLMs with human values Ouyang et al. (2022). Around the same time, Anthropic also released its own research on RLHF, demonstrating its effectiveness in guiding model behavior Bai et al. (2022a). These efforts treated safety (or harmlessness) as a subset of preference optimization, laying the foundation for future safety alignment research. From a high-level perspective, subsequent alignment techniques can be approximately categorized into the following approaches:
In-Context Learning
@@
This category of methods does not rely on model retraining; instead, inspired by Chain-of-Thought (CoT) reasoning and GPT-3, researchers have designed either hard prompts or soft prompts to guide the model toward producing helpful and harmless outputs Wei et al. (2022); Brown et al. (2020). Examples include the official system prompt in Llama2 and soft prompts optimized via P-tuning, which embed safety reasoning signals that may not be readable by humans Touvron et al. (2023); Xie et al. (2023). However, in-context learning has several limitations: (1) It requires careful manual design and optimization, making automation and scalability difficult. (2) It has limited generalization ability, struggling to handle long-tail scenarios where prompts may not be well-defined. (3) It is highly sensitive to prompt design; small variations in the prompt can lead to drastically different outputs, making the method vulnerable to jailbreak attacks. (4) Since in-context learning does not modify the model’s weights, it cannot permanently alter its underlying behavior.
Imitation Learning
@@
This approach primarily relies on supervised fine-tuning (SFT) to train models using carefully curated aligned datasets Ouyang et al. (2022)Zhou et al. (2024a) introduced the Superficial Alignment Hypothesis, suggesting that alignment may merely adjust the model’s output distribution to be more interaction-friendly rather than fundamentally changing its reasoning capabilities. They demonstrated that with only about 1k high-quality training examples, a model could achieve performance comparable to GPT-4. However, this paper focused predominantly on helpfulness, and its training data contained only 13 safety-related samples, which led to poor overall safety performance.
Reinforcement Learning
@@
Reinforcement Learning with Human Feedback (RLHF) improves model alignment by optimizing the policy (parameters) using reinforcement learning. The core idea is to train the model to generate better responses by maximizing a reward function, which is defined by a reward model trained on human feedback. The reward model assigns scores to different responses, providing a structured signal to guide the optimization process. To prevent the model from diverging too far from its original behavior, RLHF typically employs proximal policy optimization (PPO), which introduces a KL divergence constraint between the logits of the original and updated policies Christiano et al. (2017). This ensures that while the model learns to produce more aligned outputs, it does not lose fluency or develop unintended artifacts. RLHF has significantly improved both helpfulness and harmlessness, establishing itself as a foundational technique in alignment research Ouyang et al. (2022); Bai et al. (2022a). Despite its effectiveness, RLHF comes with significant challenges. The dependence on human feedback makes it highly labor-intensive, as continuous human involvement is required to annotate responses and update the reward model. To reduce reliance on human labor and improve scalability, researchers have proposed Reinforcement Learning with AI Feedback (RLAIF) as an alternative. For example, Anthropic incorporates an AI-generated feedback mechanism, where the model itself evaluates responses, reducing the need for human intervention while still refining behavior and mitigating harmful outputs Bai et al. (2022b). However, both RLHF and RLAF are resources intensive, as their training typically involves the following components: (1) a reference model, (2) a policy model (an adapted version of the base model being optimized), and (3) a reward model trained to assess response quality. In some implementations, additional models may even be required, further increasing the memory and computational burden Yao et al. (2023).
Contrastive Learning
@@
Researchers have explored contrastive learning as a more efficient alternative to reduce the complexity and resource demands of RL-based approaches. Direct Preference Optimization (DPO) introduced the key insight that large language models inherently act as implicit reward models Rafailov et al. (2024). By leveraging the preference dataset, a model can directly learn preferences without requiring an explicit reinforcement learning loop and a reward model. A similar approach is proposed in Liu et al. (2023), where contrastive signals are embedded in datasets to steer models toward desired behaviors. This category of methods has notable advantages: (1) it significantly reduces memory costs since optimization requires loading at most two models at a time, and with logit caching, only one model may be sufficient; (2) it eliminates the need for an explicit reward model, simplifying the alignment process. However, these methods also come with challenges: the performance is highly dependent on the quality of the preference dataset.
Conditional Learning
@@
This approach is conceptually similar to in-context learning but differs in that it explicitly optimizes the model to recognize specific triggers, ensuring that desired behaviors are always generated when these triggers are present. The key idea is to induce the model to produce desired behavior rather than removing undesired outputs. However, this approach has significant vulnerabilities when confronted with jailbreak attacks and thus is rarely used as a standalone alignment technique Korbak et al. (2023).
@@ 2.3 Safety in Existing LLMs is Still Brittle
Although various general alignment methods have been proposed, and safety alignment has been improved to some extent, treating safety merely as a subset of human preference overlooks its unique challenges. As a result, current alignment techniques remain vulnerable to adversarial attacks. In literature, adversarial attacks on LLMs can generally be classified into three types: (1) Jailbreak Attacks: Attackers exploit techniques such as role-playing or suffix injections to bypass safety guardrails and manipulate the model into generating harmful content. Studies show that these methods can effectively evade existing alignment mechanisms Zou et al. (2023). This vulnerability extends beyond open-source models—even state-of-the-art systems like the GPT-4 series struggle to block harmful outputs in complex, nested scenarios consistently Li et al. (2023). (2) Finetuning Attacks: Even unintentional finetuning can weaken a model’s safety mechanisms. A model trained with safety alignment may gradually lose its safeguards when adapted to downstream tasks via domain-specific finetuning, even if the dataset itself is benign. This phenomenon has been observed in both open-source and proprietary models Qi et al. (2023). (3) Decoding Attacks: Safety-aligned models may still produce harmful content under certain decoding settings, such as modifications to Top-P, Top-K, or Temperature Huang et al. (2023). These variations may break built-in safeguards, leading to outputs that would otherwise be restricted under default configurations. These attack vectors underscore a critical issue: existing safety alignment methods lack robustness and, in many cases, remain highly brittle, especially in novel or adversarial conditions.
@@ 2.4 How to Implement Robust Safety in LLMs?
Recent studies have highlighted that existing alignment methods often achieve safety at a superficial level. Wei et al. (2024) identified safety-critical parameters in LLMs and found that removing them catastrophically degrades safety performance while leaving utility performance unaffected. However, their findings also revealed that merely retaining these safety-critical parameters does not preserve safety under finetuning attacks. In contrast, Li and Kim (2024) demonstrated that the atomic functional unit for safety in LLMs resides at the neuron level and successfully mitigated finetuning attacks by freezing updates to these safety-critical components. Their study further showed that aligned models remain vulnerable to finetuning attacks because key attributes, such as utility, can be achieved by repurposing neurons originally responsible for other functions, such as safety. Additionally, this research examined how alignment influences model behavior in safety-critical contexts and observed that, at its core, this effect could be framed as an implicit safety-related binary classification task. To resolve the superficiality issue above, they further propose that alignment should enable models to choose the correct safety-aware reasoning direction (either to refuse or fulfill) at each generation step, ensuring safety throughout the entire response. However, their work did not propose specific methods for implementing this deeper safety mechanism in practice.
Qi et al. (2024) have also examined the shallow alignment in existing LLMs and found that this issue often stems from alignment disproportionately affecting early-generated token distribution. This creates optimization shortcuts where models rely on superficial decision patterns, leading them toward local optima that fail to generalize to more complex safety challenges. To mitigate this, they introduced a data augmentation strategy designed to expose models to more nuanced scenarios where an initially harmful response later transitions into a safe refusal. Similarly, Yuan et al. (2024) have adopted more aggressive data construction rules, aiming to add more variety of training examples. However, while these methods increase the diversity of training examples, they do not fundamentally address the root problem. All of these highlight a critical issue: Existing alignment techniques lack effective and robust mechanisms to handle complex and nuanced harmful reasoning patterns. In this context, this survey paper acknowledges the hypothesis from Li and Kim (2024), and believes that a robust safety alignment should teach the model to select and maintain the correct safety reasoning direction throughout the entire text generation process.
@@ 3 Spurious Biases and their Impact on Generalizability
Deep neural networks tend to learn and rely on correlations between partly predictive spurious features that are causally unrelated to ground truth labels in the training data. For example, assume one wants to train a deep neural network to be able to correctly classify pictures of animals as Cows or Camels Arjovsky et al. (2019). Due to selection bias, most samples that have Cows in them are present in green backgrounds in the training set while most Camels are present in brown backgrounds. In such a setting, deep networks are shown to learn the correlation between the background color and the ground truth labels. Such correlations are referred to as spurious correlations. In practice, deep networks often prefer spurious correlations over correlations between fully predictive, general features (features of Cows or Camels) and ground truth labels. The learning of and reliance on these correlations is undesirable because these features may disappear or become correlated with a different label or task during testing, causing these networks to malfunction.
@@ 3.1 Why do Deep Neural Networks Learn and Rely on Spurious Correlations?
Deep networks learn and rely on spurious correlations due to a preference for simpler features over those that are more complex in nature Geirhos et al. (2020); Kirichenko et al. (2023). Shah et al. (2020) show that such simplicity bias is extreme in practice. They consider a binary classification task, where every sample of each class contains two sets of features. One of these features is simpler than the other. They show that when a network is trained on this task, the network will fully ignore the more complex feature and rely only on the simpler feature when making predictions. In settings where the simpler feature does not exist in all samples, deep networks learn both sets of features but exhibit strong reliance on the simpler feature Kirichenko et al. (2023). All existing works that study spurious correlations generally assume the same set-up, where spurious features are only partly predictive of the task while general, invariant features exist in every sample within their respective class.
@@ 3.2 Mitigating Spurious Correlations: Existing Practice
Existing solutions that enable a network to mitigate spurious correlations operate under the implicit assumption that a network trained using Empirical Risk Minimization (ERM) Vapnik (1998) will learn and rely on spurious correlations due to a preference for simpler features. Based on this assumption, promising solutions generally fall into the following categories:
@@
The degree to which a trained network relies on spurious correlations depends on various factors. Of these factors, the most extensively studied is the proportion of samples within the train set that contain the spurious feature. The greater the proportion of samples containing the spurious feature, the greater the reliance on spurious correlations. To reduce the proportion of samples containing spurious features, existing works aim to either up-weight samples that do not contain spurious features, down-weight samples that contain spurious features, or remove samples containing spurious features. Most works that attain state-of-the-art results on popular benchmarks rely on the availability of sample-environment membership information. Liu et al. (2021) up-weight samples that do not contain spurious features while Yang et al. (2024) down-weight samples containing spurious features in conjunction with a similar up-weighting step. Kirichenko et al. (2023); Deng et al. (2023) simply balance the number of samples belonging to each environment when proposing mitigation strategies. However, they make use of the assumption that environments that are overrepresented force networks to rely on spurious correlations. Attaining such sample-environment information is expensive due to the need for human intervention and annotation. To overcome this problem, some works aim to infer such sample-wise environment labels. Liu et al. (2021) train a network with heavy regularization to identify samples with and without spurious features based on whether these samples were correctly classified during training. Ahmed et al. (2021) aim to maximize the Invariant Risk Minimization penalty (IRM) Arjovsky et al. (2019) during training to obtain environment-labels. Zhang et al. (2022) cluster a biased network’s representations to obtain these labels. Pezeshki et al. (2024) attain these labels by utilizing a twin-network setting where networks are encouraged to learn environmental cues, thereby aiding in sample-environment discovery.
@@
These works either align the representation of samples within a class that contains spurious features and those that do not, or simply block parts of a network’s representation that encodes spurious information. Ahmed et al. (2021) aim to align the predicted distributions for samples belonging to the same class but different environments using a KL-divergence term in the optimization function. Zhang et al. (2022) make use of a contrastive loss function which brings representations of samples within the same class but different environments closer while distancing representations of samples belonging to the same environment but different classes. Gandelsman et al. (2024) identify the role of individual attention heads in CLIP-ViT and remove those heads associated with spurious cues.
@@
Sagawa et al. (2020) optimize a network using an objective that minimizes the risk for the group of samples belonging to the environment with the maximum risk within a class.
@@
Kirichenko et al. (2023) re-train the last layer of a trained (biased) network on a dataset where the proportion of samples containing the spurious feature is significantly lower than the original training set. Moayeri et al. (2023) follow similar retraining, where they fine-tune a trained network on a small dataset with minimal spurious features, where such a set is obtained using human supervision.
Figure 1: Excluding only a handful of training samples with spurious features and hard core features mitigates spurious correlations. This is indicated by high Worst Group Accuracies (Female test samples with glasses.) Excluding up to 97% of all training samples with spurious features and easy core features shows no improvements in worst group accuracy. This figure is excerpted from Mulchandani and Kim (2025).

### v2

@@ Abstract
keywords:
Trustworthy AI, AI Safety, Spurious correlation, Privacy
@@ 2 Safety Alignment in LLMs
In Large Language Models, alignment aims to teach models human-desired behaviors and remove undesired behaviors. Safety alignment has often been treated as a subset of broader alignment challenges, with a primary focus on safety [1]. In this context, the goal of safety alignment is preventing LLMs from generating toxic or harmful content and simultaneously considers security problems in adversarial scenarios, such as jailbreak attempts [2].
@@ 2.1 Why is Safety in LLMs Important?
With the release of AI services such as ChatGPT, Claude, and Gemini, AI-powered applications have become increasingly integrated into our daily lives, spanning fields like healthcare, finance, education, transportation, and even military applications [3, 4]. However, this rapid adoption also raises serious concerns regarding AI misuse. For instance, in 2023, the first suspected case of AI-assisted suicide was reported [5], and AI-generated misinformation has been widely disseminated online, potentially manipulating public opinion [6]. These incidents compel us to critically examine how to prevent AI from being misused in ways that harm society.
This challenge has become even more pressing with the rise of open-sourced large language models such as Llama and Deepseek families [7, 8, 9], which allow individuals to control and finetune LLMs directly. As a result, the risks associated with AI misuse are expanding exponentially while government regulatory measures struggle to keep pace. Given these developments, AI safety has become a pressing need in the current research community.
@@ 2.2 How to Implement Safety in LLMs?
Upon the release of GPT-3 in 2020, we witnessed LLM’s remarkable language generation capabilities. However, concerns regarding bias, toxic content, and hallucinations also emerged, indicating that the model was still not ready for public release [10]. Later, in 2021, Anthropic introduced the HHH principle—Helpful, Honest, and Harmless—as a key principle of a truly beneficial AI assistant [11]. This concept set a foundational standard for AI assistants, clearly indicating that safety is an important objective of LLMs.
@@
The launch of ChatGPT in late 2022 marked a turning point, as it was the first large-scale exposure of LLMs to the public, allowing people to experience an AI assistant that felt genuinely helpful and capable of human-like reasoning. This breakthrough was largely built upon the techniques outlined in the InstructGPT, which introduced Supervised Fine-Tuning (SFT) and Reinforcement Learning with Human Feedback (RLHF) as key methods for aligning LLMs with human values [12]. Around the same time, Anthropic also released its own research on RLHF, demonstrating its effectiveness in guiding model behavior [13]. These efforts treated safety (or harmlessness) as a subset of preference optimization, laying the foundation for future safety alignment research. From a high-level perspective, subsequent alignment techniques can be approximately categorized into the following approaches:
2.2.1 In-Context Learning
@@
This category of methods does not rely on model retraining; instead, inspired by Chain-of-Thought (CoT) reasoning and GPT-3, researchers have designed either hard prompts or soft prompts to guide the model toward producing helpful and harmless outputs [14, 10]. Examples include the official system prompt in Llama2 and soft prompts optimized via P-tuning, which embed safety reasoning signals that may not be readable by humans [7, 15]. However, in-context learning has several limitations: (1) It requires careful manual design and optimization, making automation and scalability difficult. (2) It has limited generalization ability, struggling to handle long-tail scenarios where prompts may not be well-defined. (3) It is highly sensitive to prompt design; small variations in the prompt can lead to drastically different outputs, making the method vulnerable to jailbreak attacks. (4) Since in-context learning does not modify the model’s weights, it cannot permanently alter its underlying behavior.
2.2.2 Imitation Learning
@@
This approach primarily relies on supervised fine-tuning (SFT) to train models using carefully curated aligned datasets [12][16] introduced the Superficial Alignment Hypothesis, suggesting that alignment may merely adjust the model’s output distribution to be more interaction-friendly rather than fundamentally changing its reasoning capabilities. They demonstrated that with only about 1k high-quality training examples, a model could achieve performance comparable to GPT-4. However, this paper focused predominantly on helpfulness, and its training data contained only 13 safety-related samples, which led to poor overall safety performance.
2.2.3 Reinforcement Learning
@@
Reinforcement Learning with Human Feedback (RLHF) improves model alignment by optimizing the policy (parameters) using reinforcement learning. The core idea is to train the model to generate better responses by maximizing a reward function, which is defined by a reward model trained on human feedback. The reward model assigns scores to different responses, providing a structured signal to guide the optimization process. To prevent the model from diverging too far from its original behavior, RLHF typically employs proximal policy optimization (PPO), which introduces a KL divergence constraint between the logits of the original and updated policies [17]. This ensures that while the model learns to produce more aligned outputs, it does not lose fluency or develop unintended artifacts. RLHF has significantly improved both helpfulness and harmlessness, establishing itself as a foundational technique in alignment research [12, 13]. Despite its effectiveness, RLHF comes with significant challenges. The dependence on human feedback makes it highly labor-intensive, as continuous human involvement is required to annotate responses and update the reward model. To reduce reliance on human labor and improve scalability, researchers have proposed Reinforcement Learning with AI Feedback (RLAIF) as an alternative. For example, Anthropic incorporates an AI-generated feedback mechanism, where the model itself evaluates responses, reducing the need for human intervention while still refining behavior and mitigating harmful outputs [18]. However, both RLHF and RLAF are resources intensive, as their training typically involves the following components: (1) a reference model, (2) a policy model (an adapted version of the base model being optimized), and (3) a reward model trained to assess response quality. In some implementations, additional models may even be required, further increasing the memory and computational burden [19].
2.2.4 Contrastive Preference Modeling
@@
Researchers have explored the use of contrastive preference signals as a more efficient alternative to reduce the complexity and resource demands of RL-based approaches. Direct Preference Optimization (DPO) introduced the key insight that large language models inherently act as implicit reward models [20]. By leveraging the preference dataset, a model can directly learn preferences without requiring an explicit reinforcement learning loop and a reward model. A similar approach is proposed in [21], where contrastive signals are embedded in datasets to steer models toward desired behaviors. This category of methods has notable advantages: (1) it significantly reduces memory costs since optimization requires loading at most two models at a time, and with logit caching, only one model may be sufficient; (2) it eliminates the need for an explicit reward model, simplifying the alignment process. However, these methods also come with challenges: the performance is highly dependent on the quality of the preference dataset.
2.2.5 Conditional Learning
@@
This approach is conceptually similar to in-context learning but differs in that it explicitly optimizes the model to recognize specific triggers, ensuring that desired behaviors are always generated when these triggers are present. The key idea is to induce the model to produce desired behavior rather than removing undesired outputs. However, this approach has significant vulnerabilities when confronted with jailbreak attacks and thus is rarely used as a standalone alignment technique [22].
@@ 2.3 Safety in Existing LLMs is Still Brittle
Although various general alignment methods have been proposed, and safety alignment has been improved to some extent, treating safety merely as a subset of human preference overlooks its unique challenges. As a result, current alignment techniques remain vulnerable to adversarial attacks. In literature, adversarial attacks on LLMs can generally be classified into three types: (1) Jailbreak Attacks: Attackers exploit techniques such as role-playing or suffix injections to bypass safety guardrails and manipulate the model into generating harmful content. Studies show that these methods can effectively evade existing alignment mechanisms [23]. This vulnerability extends beyond open-source models—even state-of-the-art systems like the GPT-4 series struggle to block harmful outputs in complex, nested scenarios consistently [24]. (2) Finetuning Attacks: Even unintentional finetuning can weaken a model’s safety mechanisms. A model trained with safety alignment may gradually lose its safeguards when adapted to downstream tasks via domain-specific finetuning, even if the dataset itself is benign. This phenomenon has been observed in both open-source and proprietary models [25]. (3) Decoding Attacks: Safety-aligned models may still produce harmful content under certain decoding settings, such as modifications to Top-P, Top-K, or Temperature [26]. These variations may break built-in safeguards, leading to outputs that would otherwise be restricted under default configurations. These attack vectors underscore a critical issue: existing safety alignment methods lack robustness and, in many cases, remain highly brittle, especially in novel or adversarial conditions.
@@ 2.4 How to Implement Robust Safety in LLMs?
Recent studies have highlighted that existing alignment methods often achieve safety at a superficial level. [27] identified safety-critical parameters in LLMs and found that removing them catastrophically degrades safety performance while leaving utility performance unaffected. However, their findings also revealed that merely retaining these safety-critical parameters does not preserve safety under finetuning attacks. In contrast, [1] demonstrated that the atomic functional unit for safety in LLMs resides at the neuron level and successfully mitigated finetuning attacks by freezing updates to these safety-critical components. Their study further showed that aligned models remain vulnerable to finetuning attacks because key attributes, such as utility, can be achieved by repurposing neurons originally responsible for other functions, such as safety. Additionally, this research examined how alignment influences model behavior in safety-critical contexts and observed that, at its core, this effect could be framed as an implicit safety-related binary classification task. To resolve the superficiality issue above, they further propose that alignment should enable models to choose the correct safety-aware reasoning direction (either to refuse or fulfill) at each generation step, ensuring safety throughout the entire response. However, their work did not propose specific methods for implementing this deeper safety mechanism in practice.
[2] have also examined the shallow alignment in existing LLMs and found that this issue often stems from alignment disproportionately affecting early-generated token distribution. This creates optimization shortcuts where models rely on superficial decision patterns, leading them toward local optima that fail to generalize to more complex safety challenges. To mitigate this, they introduced a data augmentation strategy designed to expose models to more nuanced scenarios where an initially harmful response later transitions into a safe refusal. Similarly, [28] have adopted more aggressive data construction rules, aiming to add more variety of training examples. However, while these methods increase the diversity of training examples, they do not fundamentally address the root problem. All of these highlight a critical issue: Existing alignment techniques lack effective and robust mechanisms to handle complex and nuanced harmful reasoning patterns. In this context, this survey paper acknowledges the hypothesis from [1], and believes that a robust safety alignment should teach the model to select and maintain the correct safety reasoning direction throughout the entire text generation process. This perspective is aligned with recent work in [29], which not only supports this view but also introduces practical techniques to enforce such reasoning consistency.
@@ 3 Spurious Biases and their Impact on Generalizability
Deep neural networks tend to learn and rely on correlations between partly predictive spurious features that are causally unrelated to ground truth labels in the training data. For example, assume one wants to train a deep neural network to be able to correctly classify pictures of animals as Cows or Camels [30]. Due to selection bias, most samples that have Cows in them are present in green backgrounds in the training set while most Camels are present in brown backgrounds. In such a setting, deep networks are shown to learn the correlation between the background color and the ground truth labels. Such correlations are referred to as spurious correlations. In practice, deep networks often prefer spurious correlations over correlations between fully predictive, general features (features of Cows or Camels) and ground truth labels. The learning of and reliance on these correlations is undesirable because these features may disappear or become correlated with a different label or task during testing, causing these networks to malfunction.
@@ 3.1 Why do Deep Neural Networks Learn and Rely on Spurious Correlations?
Deep networks learn and rely on spurious correlations due to a preference for simpler features over those that are more complex in nature [31, 32]. [33] show that such simplicity bias is extreme in practice. They consider a binary classification task, where every sample of each class contains two sets of features. One of these features is simpler than the other. They show that when a network is trained on this task, the network will fully ignore the more complex feature and rely only on the simpler feature when making predictions. In settings where the simpler feature does not exist in all samples, deep networks learn both sets of features but exhibit strong reliance on the simpler feature [32]. All existing works that study spurious correlations generally assume the same set-up, where spurious features are only partly predictive of the task while general, invariant features exist in every sample within their respective class.
@@ 3.2 Mitigating Spurious Correlations: Existing Practice
Existing solutions that enable a network to mitigate spurious correlations operate under the implicit assumption that a network trained using Empirical Risk Minimization (ERM) [34] will learn and rely on spurious correlations due to a preference for simpler features. Based on this assumption, promising solutions generally fall into the following categories:
@@
The degree to which a trained network relies on spurious correlations depends on various factors. Of these factors, the most extensively studied is the proportion of samples within the train set that contain the spurious feature. The greater the proportion of samples containing the spurious feature, the greater the reliance on spurious correlations. To reduce the proportion of samples containing spurious features, existing works aim to either up-weight samples that do not contain spurious features, down-weight samples that contain spurious features, or remove samples containing spurious features. Most works that attain state-of-the-art results on popular benchmarks rely on the availability of sample-environment membership information. [35] up-weight samples that do not contain spurious features while [36] down-weight samples containing spurious features in conjunction with a similar up-weighting step. [32, 37] simply balance the number of samples belonging to each environment when proposing mitigation strategies. However, they make use of the assumption that environments that are overrepresented force networks to rely on spurious correlations. Attaining such sample-environment information is expensive due to the need for human intervention and annotation. To overcome this problem, some works aim to infer such sample-wise environment labels. [35] train a network with heavy regularization to identify samples with and without spurious features based on whether these samples were correctly classified during training. [38] aim to maximize the Invariant Risk Minimization penalty (IRM) [30] during training to obtain environment-labels. [39] cluster a biased network’s representations to obtain these labels. [40] attain these labels by utilizing a twin-network setting where networks are encouraged to learn environmental cues, thereby aiding in sample-environment discovery.
@@
These works either align the representation of samples within a class that contains spurious features and those that do not, or simply block parts of a network’s representation that encodes spurious information. [38] aim to align the predicted distributions for samples belonging to the same class but different environments using a KL-divergence term in the optimization function. [39] make use of a contrastive loss function which brings representations of samples within the same class but different environments closer while distancing representations of samples belonging to the same environment but different classes. [41] identify the role of individual attention heads in CLIP-ViT and remove those heads associated with spurious cues.
@@
[42] optimize a network using an objective that minimizes the risk for the group of samples belonging to the environment with the maximum risk within a class.
@@
[32] re-train the last layer of a trained (biased) network on a dataset where the proportion of samples containing the spurious feature is significantly lower than the original training set. [43] follow similar retraining, where they fine-tune a trained network on a small dataset with minimal spurious features, where such a set is obtained using human supervision.
Figure 1: Excluding only a handful of training samples with spurious features and hard core features mitigates spurious correlations. This is indicated by high Worst Group Accuracies (Female test samples with glasses.) Excluding up to 97% of all training samples with spurious features and easy core features shows no improvements in worst group accuracy. This figure is excerpted from [44].